# Multimodal Conversation Structure Understanding

### v2

@@ A.1 Annotation guidelines
The section outlines the conceptual framework we use to annotate conversational roles and threads in multi-party conversation. [^7] Written primarily by AH and KC, with input and guidance from other co-authors. We draw on such traditions as sociolinguistics, linguistic anthropology, and television studies, especially frameworks developed by Clark and Murphy [44, 45], Sacks and Schegloff [60], and Goodwin [43], who, taken together, lay the foundation of conversation analysis and provide us with the basis of taxonomy on which our role attribution task is based; Goffman [1], on the face-to-face, social nature of dialogic interactions; and finally, McKee [61], who analyzes on TV dialogues in the familiar set of vocabulary (Austinian speech acts, for instance).
Our annotation scheme is the result of several iterations, beginning with a pilot round and subsequent discussions with annotators trained in rhetorics, cognitive science, film studies, and related disciplines. Based on observations from the pilot, we refine the definitions of utterance boundaries, reply-to structure, and role attribution to better reflect how participants—or characters on screen—engage in face-to-face dialogues in TVQA data. We hope the annotation guidelines presented below will bring out the nuances of the annotation tasks at hand and their intellectual stakes, as well as provide the theoretical scaffolding that informs our design choices.
In any dialogic interaction, conversational participants manage and direct attention towards each other, during which they also inhibit various roles (speakers, addressees, etc.). For Goffman [1] in particular, conversations are more than merely linguistic exchanges but ritualized social encounters, regulated by norms that govern such phenomena as turn-taking and floor-claiming. These norms extend beyond linguistic structures alone and are often mediated through non-verbal and non-linguistic cues (gaze, posture, etc.).
@@
This draws on McKee’s notion of speech acts, as characters who initiate dialogue to perform an action, and Goffman’s idea that conversation involves the distribution of attention among its participants. Characters who get to start a conversation and main the floor can be more powerful or otherwise interesting.
@@
Building on Goffman’s notion of adjacency pairs [60] and Clark’s ratified vs. non-ratified participation, participants who are present but not addressed have the potential of illuminating social hierarchies, inclusion/exclusion patterns, and attentional structures within multi-party conversations.
The annotation guidelines are organized as follows: §A.1.1: Definitions; §A.1.2: Annotation procedure; §A.1.3: Examples; §A.1.4: Annotation quality; §A.1.5: Limitations.
A.1.1 Definitions
This is grounded primarily in [1, 61]:
@@ A.1.1 Definitions
An utterance is a single unit of spoken dialogue, roughly equivalent to a dialogue line. It is a bounded communicative act performed by a speaker within an interaction driven by the need for a plot, of the speaker, or other characters in the context of a TV show. Unlike a statement, which is defined by its propositional content and logical truth conditions, an utterance is defined by its timing, delivery, and interactional role. Utterances need not be syntactically complete sentences, but as moves that accomplish social action. A particular type of utterance is the utterance of interest (UOI), which is the current utterance being annotated. For ease of annotation, we present individual utterances as one sentence transcribed by Whisper [36]. They don’t have to be complete sentences, and they don’t have to be a completed turn in and of itself. This allows us to annotate conversational roles at a reasonable granularity: different lines in the same turn can have different role attributions.
@@
The reply-to relation connects an utterance of interest to its most relevant preceding utterance, also known as a parent utterance, representing the line to which it responds most directly. This is more prevalent in NLP literature [62], where the task is formalized as a link from a child utterance to its parent, and the relationship is modeled as directed graphs. The default parent utterance is the immediately preceding line; if there is no logical antecedent utterance, the UOI is the beginning of a new thread (more below). Since we present each utterance at the sentence level, a special form of reply-to is that of a continuation: if the speaker is still in the midst of their turn, and the UOI merely continues the previous line, then the previous line is the natural and logical parent utterance. In this view, we might see TV dialogues as linear sequences, where one line necessarily triggers and informs the next, until the current thread runs its course.
@@
For more information, see the Appendix of [40], including specific tests and examples.
@@
The first role is the speaker, the animator, the source of the utterance at time $t$. Following Goodwin’s conversational analysis [43] and Clark’s role taxonomy [44, 45], we treat conversational roles as those observably projected at $t$: a speaker’s alignment towards others (signaled through gaze, body orientation, lexical choice) casts other participants into roles. These roles are independent of whether the projected utterance ultimately hears, understands, or replies.
@@
Role addressed ratified known
Addressee $+$ $+$ $+$
Side-participants -- $+$ $+$
Bystanders -- -- $\pm$
@@
  • Side-participant: a participant who is ratified, perceptually engaged, but not explicitly addressed. This term is not really used in Goffman, but we prefer it because it has been used in previous work related to ours (e.g., [63]).
@@ A.1.2 Annotation procedure
Figure 4: Annotation enhancements for edge cases and workflow efficiency. Annotators can select characters and special participant tokens for non-standard utterances. Checkboxes allow them to mark lines as continuations (same role as previous) or non-dialogic (extra-diagetic or monologue), with the goal to minimize redundancy and improve overall annotation consistency.
@@
  • Extra-diagetic (checkbox): to indicate narrators, “previously on the show”, or other kinds of dialogue lines that are not part of any actual conversation
@@ A.1.3 Examples
Here are some edge cases:
@@
In this scene from House M. D. (season 4, episode 16, segment 2, clip 10), a roomful of doctors speak amongst themselves while one doctor in particular, James Wilson, sits on the side and remains largely unspeaking and unshown. Though the camera is rarely on him and the other doctors address each other rather than him, he qualifies as a side-participant due to being in the same space and scene as the other interlocutors.
@@
A.1.4 Annotation quality
We report inter-annotator agreement in Table 7 based on 10–12 sampled clips (not part of the final dataset). Agreement is computed as the average of all pairwise comparisons among the four primary annotators, using the same set of evaluation metrics for experiments, and we observe a clear improvement in the main round of annotation.
Table 7: Annotation quality.
[…]
A.1.5 Limitations
@@
A.2 Additional experimental details
class ConversationalRoles(BaseModel):
line_index: int
reply_to: int
speaker: str
addressees: list[str]
side_participants: list[str]
class ClipRoles(BaseModel):
clip_roles: list[ConversationalRoles]
Figure 5: Pydantic schema.
We do not change the default temperature of the LLMs that we evaluate. We use the Pydantic schema for structured output (see Fig. 5). Our system instructions for Gemini 2.0 Flash, similar to those used for other models we consider, are in Fig. 6. This represents a sunccinct summary of our annotation guidelines. For our LLaMA 4 Scout experiments, we run the model in full precision on four H100 GPUs served by vLLM [64], which takes around 3 hours to complete. Qwen 2.5-Omni 7B inference is performed with full precision with Flash Attention 2 enabled [65] on four L40S GPUs, which takes around 7 hours. For Qwen 2.5-Omni 7B LoRA SFT, we follow the instructions and configurations provided on https://github.com/hiyouga/LLaMA-Factory/pull/7537.

### v3

@@ Appendix F Annotation guidelines
The section outlines the conceptual framework we use to annotate conversational roles and threads in multi-party conversation. [^8] Written primarily by AH and KC, with input and guidance from other co-authors. We draw on such traditions as sociolinguistics, linguistic anthropology, and television studies, especially frameworks developed by Clark and Carlson (1982); Clark and Schaefer (1987)Sacks et al. (1974), and Goodwin (1981), who, taken together, lay the foundation of conversation analysis and provide us with the basis of taxonomy on which our role attribution task is based; Goffman (1981), on the face-to-face, social nature of dialogic interactions; and finally, McKee (2016), who analyzes TV dialogues in the familiar set of vocabulary (Austinian speech acts, for instance).
Our annotation scheme is the result of several iterations, beginning with a pilot round and subsequent discussions with annotators trained in rhetoric, cognitive science, film studies, and related disciplines. Based on observations from the pilot, we refine the definitions of utterance boundaries, reply-to structure, and role attribution to better reflect how participants—or characters on screen—engage in face-to-face dialogues in TVQA data. [^9] This data is obtained from the original authors under TVQA Dataset Download Agreement; see https://docs.google.com/forms/d/e/1FAIpQLSftzvAtUSnsmuwpKk5PU9bS7lk4FOu_v9Y2vq2nGjRXKoh6tw/viewform We hope the annotation guidelines presented below will bring out the nuances of the annotation tasks at hand and their intellectual stakes, as well as provide the theoretical scaffolding that informs our design choices.
In any dialogic interaction, conversational participants manage and direct attention towards each other, during which they also inhabit various roles (speakers, addressees, etc.). For Goffman (1981) in particular, conversations are more than merely linguistic exchanges but ritualized social encounters, regulated by norms that govern such phenomena as turn-taking and floor-claiming. These norms extend beyond linguistic structures alone and are often mediated through non-verbal and non-linguistic cues (gaze, posture, etc.).
@@
This draws on McKee’s notion of speech acts, as characters who initiate dialogue to perform an action, and Goffman’s idea that conversation involves the distribution of attention among its participants. Characters who get to start a conversation and maintain the floor can be more powerful or otherwise interesting.
@@
Building on Goffman’s notion of adjacency pairs Sacks et al. (1974) and Clark’s ratified vs. non-ratified participation, participants who are present but not addressed have the potential of illuminating social hierarchies, inclusion/exclusion patterns, and attentional structures within multi-party conversations.
The annotation guidelines are organized as follows: §F.1: Definitions; §F.2: Annotation procedure; §F.3: Examples; §F.5: Limitations.
F.1 Definitions
This is grounded primarily in Goffman (1981); McKee (2016):
@@ F.1 Definitions
An utterance is a single unit of spoken dialogue, roughly equivalent to a dialogue line. It is a bounded communicative act performed by a speaker within an interaction driven by the need for a plot, of the speaker, or other characters in the context of a TV show. Unlike a statement, which is defined by its propositional content and logical truth conditions, an utterance is defined by its timing, delivery, and interactional role. Utterances need not be syntactically complete sentences, but as moves that accomplish social action. A particular type of utterance is the utterance of interest (UOI), which is the current utterance being annotated. For ease of annotation, we present individual utterances as one sentence transcribed by Whisper Radford et al. (2022). They don’t have to be complete sentences, and they don’t have to be a completed turn in and of itself. This allows us to annotate conversational roles at a reasonable granularity: different lines in the same turn can have different role attributions.
@@
The reply-to relation connects an utterance of interest to its most relevant preceding utterance, also known as a parent utterance, representing the line to which it responds most directly. This is more prevalent in NLP literature Zhu et al. (2021), where the task is formalized as a link from a child utterance to its parent, and the relationship is modeled as directed graphs. The default parent utterance is the immediately preceding line; if there is no logical antecedent utterance, the UOI is the beginning of a new thread (more below). Since we present each utterance at the sentence level, a special form of reply-to is that of a continuation: if the speaker is still in the midst of their turn, and the UOI merely continues the previous line, then the previous line is the natural and logical parent utterance. In this view, we might see TV dialogues as linear sequences, where one line necessarily triggers and informs the next, until the current thread runs its course.
@@
The first role is the speaker, the animator, the source of the utterance at time $t$. Following Goodwin’s conversational analysis Goodwin (1981) and Clark’s role taxonomy Clark and Carlson (1982); Clark and Schaefer (1987), we treat conversational roles as those observably projected at $t$: a speaker’s alignment towards others (signaled through gaze, body orientation, lexical choice) casts other participants into roles. These roles are independent of whether the projected utterance ultimately hears, understands, or replies.
Role addressed ratified known Addressee ++ ++ ++ Side-participants - ++ ++ Bystanders - - $\pm$
@@
  • Side-participant: a participant who is ratified, perceptually engaged, but not explicitly addressed. This term is not really used in Goffman, but we prefer it because it has been used in previous work related to ours (e.g., Lee and Marsella (2011)).
@@ F.2 Annotation procedure
Figure 8: Annotation enhancements for edge cases and workflow efficiency. Annotators can select characters and special participant tokens for non-standard utterances. Checkboxes allow them to mark lines as continuations (same role as previous) or non-dialogic (extra-diegetic or monologue), with the goal to minimize redundancy and improve overall annotation consistency.
@@
  • Extra-diegetic (checkbox): to indicate narrators, “previously on the show”, or other kinds of dialogue lines that are not part of any actual conversation
@@ F.3 Examples
F.4 Edge cases
@@
In this scene from House M. D. (season 4, episode 16, segment 2, clip 10), a roomful of doctors speak amongst themselves while one doctor in particular, James Wilson, sits on the side and remains largely unspeaking and unshown. Even though the camera is rarely on him and the other doctors address each other rather than him, he qualifies as a side-participant due to being in the same space and scene as the other interlocutors.
@@
F.5 Limitations